An inheritance of $24,000 is to be divided among three trusts, with the second trust received twice as much as the first trust.The three trust pay interest at the rates of 9%,10%,and 6% respectively,and return a total in interest of $2210 at the end of the first year.How much was intvested in each trust?
Answer by anuttam(1)

given x+y+z=24000 and y is equal to 2 times x then
x+2x+z=24000
Z=24000-3x
3x+z=24000 equation 1
9x+10Y+6z=221000 equation 2
multiply equation1 by -3 then add to equation 2
10y+3z=149000 equation 3
substitue y=2x and z values
20x+3(24000-3x)=149000
20x+72000-9x=149000
x=7000,y=14000,z=3000 which equals $24000 and verify that 9%,10% and 6% interest respectively on x,y,z values brings 2210 interest.